OpenStudy (anonymous):

A POLICE CAR MOVING 140KM/H CHASES A SPEEDING CAR MOVING AT 110KM/HR. HOW MUCH CLOSER THE SPEEDING CAR DOES THE POLICE CAR GET EACH MINUTES?

OpenStudy (kainui):

OK SO DUDE WHAT YOU GOT TO DO IS SUBTRACT THE SPEEDS. THINK OF IT LIKE THIS OK YOU WALK AT 5 M/H AND YOUR FRIEND WALK AT 5 M/H NOW YOU BOTH ARE WALKING THE SAME SPEED SO YOU JUST DONT CHANGE RELATIVE DISTANCE. UNDERSTAND?

OpenStudy (shane_b):

Subtract the speeds to get the difference in velocity between the two cars. \[140 km/h - 110 km/h = 30km/h\]So the police car is going 30km/h faster than the other car... The question asks how much closer does the police car get every minute. Simply convert 30km/h to km/min:\[\frac{30km}{h}*\frac{1h}{60min}=0.5\space km/min\]